OpenStudy (anonymous):

Factor 35x^2 – 2x -1

OpenStudy (kc_kennylau):

you mean factorize

OpenStudy (kc_kennylau):

(5x-1)(7x+1)

OpenStudy (kc_kennylau):

If you look at 35, you'll recognize its factors instantly: 5 and 7

OpenStudy (ranga):

35x^2 – 2x -1 35x^2 - 7x + 5x - 1 7x(5x - 1) + 1(5x - 1) (7x + 1)(5x - 1)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

using the ac method

OpenStudy (anonymous):

ax^2+ bx + c

OpenStudy (ranga):

Yes, we multiply a and c: 35 * (-1) = -35. We need to split -35 into two factors whose sum will be "b". -35 = -7 * 5 and -7 + 5 = -2 which is "b"

OpenStudy (anonymous):

just making sure it was still going to be the same answer

OpenStudy (ranga):

Yes it will be.

OpenStudy (anonymous):

Thanks Ranga